OpenStudy (anonymous):

The sum of two numbers is 52 and the difference is 6 . What are the numbers?

OpenStudy (owlcoffee):

You can solve by establishing a system of equations: \[a+b=52\] \[a-b=6\]

sammixboo (sammixboo):

x = one number y = other number x + y = 52 x - y = 6 Do you know the elimination/addition method?

sammixboo (sammixboo):

Where you add the two equations together

OpenStudy (anonymous):

no i don't

sammixboo (sammixboo):

Well adding the two equations together you do this